# Louis Serrano1  Armand Kassaï Koupaï1  Thomas X Wang1 Pierre Erbacher 2Patrick Gallinari1,3 1Sorbonne Université, CNRS, ISIR, 75005 Paris, France

### v2

@@ 4 Experiments
4.1 Datasets details
As in Kirchmeyer et al. (2022), we generate data in batches where each batch of trajectories shares the same PDE parameters. For each batch or environment, the resulting trajectories sharing the same dynamics have different initial conditions. We consider different whole factors of variations across multiple datasets and drastically increase the different number of environments compared to previous studies (Yin et al. (2022), Kirchmeyer et al. (2022)). We conduct experiments across seven datasets: five in 1D—Advection, Heat, Burgers, Wave-b, Combined—and two in 2D—Wave 2D, Vorticity. These datasets were selected to encompass different physical phenomena and test generalization under changes to various PDE terms, as described below.
Varying PDE coefficients
The changing factor is the set of coefficients $\mu$ in Equation 1. For the Burgers, Heat, and Vorticity 2D equations, the viscosity coefficient $\nu$ varies across environments. For Advection, the advection speed $\beta$ changes. In Wave-c and Wave-2D, the wave’s celerity $c$ is unique to each environment, and the damping coefficient $k$ varies across environments in Wave-2D. In the Combined equation, three coefficients $(\alpha,\beta,\gamma)$ vary, each influencing different derivative terms respectively: $-\frac{\partial u^{2}}{\partial x},+\frac{\partial^{2}u}{\partial x^{2}},-\frac{\partial^{3}u}{\partial x^{3}}$ on the right-hand side of Equation 1.
Varying boundary conditions
In this case, the varying parameter is the boundary condition $\mathcal{B}$ from Equation 2. For Wave-b, we explore two types of boundary conditions—Dirichlet and Neumann—applied independently to each boundary, resulting in four distinct environments.
Varying forcing term
The varying parameter is the forcing term $\delta$ in Equation 1. In Burgers and Heat, the forcing terms vary by the amplitude, frequency, and shift coefficients of $\delta(t,x)=\sum_{j=1}^{5}A_{j}\sin\left(\omega_{j}t+2\pi\frac{l_{j}x}{L}+\phi_{j}\right)$.
A detailed description of the datasets is provided in Appendix B, and a summary of the number of environments used during training, the number of trajectories sharing the same dynamics, and the varying PDE parameters across environments is presented in Table 2. For testing, we evaluate all methods on trajectories with new initial conditions on unseen environments. Specifically, we used 120 new environments for the 2D datasets and 12 for the 1D datasets, with each environment containing 10 trajectories.
4.2 Context adaptation from similar trajectories
Advection Heat Burgers Wave b Combined Wave 2D Vorticity 2D
CAPE 0.00941 0.223 0.213 0.978 0.00857
CODA 0.00687 0.546 0.767 1.020 0.0120 0.777 0.678
[CLS] ViT 0.140 0.136 0.116 0.971 0.0446 0.271 0.972
Zebra 0.00794 0.154 0.115 0.245 0.00965 0.207 0.119
Setting
We evaluate Zebra’s ability to perform in-context learning by leveraging example trajectories that follow the same underlying dynamics as the target. Formally, in the $n$-shot adaptation setting, we assume access to a set of $n$ context trajectories $\{{\bm{u}}_{1}^{0:m\Delta t},\ldots,{\bm{u}}_{n}^{0:m\Delta t}\}$ at inference time, all of which belong to the same dynamical system $\mathcal{F}_{\xi}$. The goal of the adaptation task is to accurately predict a future trajectory ${\bm{u}}_{*}^{\Delta t:m\Delta t}$ from a new initial condition ${\bm{u}}_{*}^{0}$, knowing that the target dynamics is shared with the provided context example trajectories. In this comparison, Zebra is the only model that performs in-context learning from these example trajectories.
Sampling
For Zebra, we employ a random sampling procedure for generating the next tokens for all datasets, setting a low temperature ($\tau=0.1$) to prioritize accuracy over diversity. Predictions are generated using a single sample under this configuration.
Baselines
We evaluate Zebra against CODA (Kirchmeyer et al., 2022) and CAPE (Takamoto et al., 2023). CODA is a meta-learning framework designed for learning parametric PDEs. It leverages common knowledge from multiple environments where trajectories from a same environment $e$ share the same PDE parameter values. CODA training performs adaptation in the parameter space by learning shared parameters across all environments and a context vector $c^{e}$ specific to each environment. At the inference stage, CODA adapts to a new environment in a one-shot manner by only tuning $c^{e}$ with several gradient steps. CAPE is not designed to perform adaptation via extra-trajectories, but instead needs the correct parameter values as input to condition a neural solver. We adapt it to our setting, by learning a context $c^{e}$ instead of using the real parameter values. During adaptation, we only tune this context $c^{e}$ via gradient updates. Additionally, we introduce a baseline based on a vision transformer (Peebles & Xie, 2023), integrating a [CLS] token that serves as a learned parameter for each environment. This token lets the model handle different dynamics, and during inference, we adapt the [CLS] vector via gradient updates, following the same approach used in CODA and CAPE. We refer to this baseline as [CLS]ViT.
Metrics
Results
As evidenced in Table 3, Zebra demonstrates strong overall performance in the one-shot adaptation setting, often surpassing baseline methods that have been trained specifically for this task. In more challenging datasets, such as Burgers, Wave-b, and the 2D cases, Zebra consistently achieves lower relative L2 errors, highlighting its capacity to model complex dynamics effectively. Notably, Zebra excels in 2D environments, outperforming both CODA and [CLS]ViT and avoiding the divergence issues encountered by CAPE. While Zebra performs comparably to CODA on simpler datasets like Advection and Combined, its overall stability and versatility across a range of scenarios, particularly in 2D settings, highlight its competitiveness. Although there is room for improvement in specific cases, such as the Heat dataset, Zebra stands out as a reliable and scalable solution for in-context adaptation for parametric PDEs, offering a robust alternative to existing gradient-based methods. We further analyze the influence of the number $n$ of context examples on the rollout performance with Zebra, as illustrated in Figure 2. While there is a general decreasing trend—indicating that more context examples tend to reduce rollout loss—there is still noticeable variance in the results. This suggests that the relationship between the number of context examples and performance is not perfectly linear. We hypothesize that this analysis would benefit from being conducted with more than a single sample per trajectory to ensure more robust estimations.
Figure 2: Influence of the number of examples. Zebra’s rollout loss for a different number of trajectory examples. The x-axis is the # of context examples and the y-axis is the Relative $L^{2}$.
4.3 Temporal conditioning
Setting
We then evaluate the temporal conditioning capabilities of Zebra, i.e. its generalization capabilities when conditioned by the initial states of the target trajectories. Formally, given a new trajectory ${\bm{u}}_{*}$, we suppose that a set of $\ell$ states ${\bm{u}}_{*}^{0:(\ell-1)\Delta t}$ is already available from the trajectory and we wish to predict the states at the following timestamps ${\bm{u}}_{*}^{\ell\Delta t:m\Delta t}$.
Baselines
We test Zebra against CODA (Kirchmeyer et al., 2022), CAPE (Takamoto et al., 2023) and MPP (McCabe et al., 2023). CODA is designed for adapting neural networks to a new environment given an extra-trajectory sampled from this environment. In this setup, we do not have access to extra-trajectories but to the first timestamps of the target trajectory. We thus modify CODA for this setting; the model is adapted with the $\ell$ first states by learning only $c^{e}$, and then starts from ${\bm{u}}^{(\ell-1)\Delta t}$ to predict the rest of the trajectory. We adapt CAPE to that setting too, as done with CODA. MPP is a vision transformer conditioned by a sequence of frames, using temporal and spatial self-attention blocks to capture spatio-temporal dependencies. MPP does not require additional adaptation at inference and can be used in zero-shot on new trajectories. It is pretrained with a fixed number of input frames. For Zebra, we employ the sampling procedure described in Section 4.2.
Table 4: Zero-shot prediction from 2 frames. Conditioning from a trajectory history with 2 frames as input. Test results in relative L2 on the trajectory. ‘–‘ indicates inference has diverged.
Advection Heat Burgers Wave b Combined Wave 2D Vorticity 2D
CAPE 0.00682 0.234 0.225 1.10 0.0125
CODA 0.00560 0.378 0.472 0.994 0.0197 0.974 0.623
MPP[2] 0.0075 0.0814 0.100 1.0393 0.0250 0.285 0.101
MPP[3] 0.919 1.0393 0.581 0.900 0.201 0.596 0.219
Zebra 0.00631 0.227 0.221 0.992 0.0084 0.201 0.0874
Results
Table 4 highlights Zebra’s strong zero-shot prediction performance using only 2 frames ($\ell=2$) as context, outperforming competing methods across a wide range of PDEs. Notably, Zebra excels on both 1D and 2D datasets, delivering consistent and robust results even in complex dynamics like Wave 2D and Vorticity 2D.
Figure 3: Zero-shot vs one-shot performance of Zebra with 2 frames.
CAPE and CODA, while competitive in some datasets, either diverge or struggle with accuracy in more challenging scenarios, particularly in 2D problems. MPP trained with two frames (MPP[2]) is overall a very strong baseline in this setting; it performs best on Heat and Burgers and obtains good results in the 2D cases. However, if we take a model that has been pretrained specifically on three frames (MPP[3]), and test it under this setting, the performance degrades drastically. In contrast, Zebra exhibits a high flexibility. It can be used with any number of frames, as long it does not exceed the maximum sequence size seen during training. Furthermore, keeping this setting with two initial frames as inputs, we expose in Figure 3 the gains we could expect on the rollout loss if we had access in addition to the input frames to an example trajectory as described in Section 4.2. We can observe that Zebra consistently improves its accuracy when prompted with an additional example. Most notably, Zebra’s behaviour goes from a random predicition on Wave b in zero-shot to more confident predictions thanks to the additional example.
4.4 Uncertainty quantification
Motivated by this observation, we investigate how varying the model’s temperature parameter $\tau$ affects its predictions; specifically in the one-shot adaptation setting described in Section 4.2. By adjusting $\tau$, we aim to assess its impact on both the accuracy and variability of the predictions. We employ three metrics to evaluate the model’s uncertainty:
Figure 4: Uncertainty quantification with Zebra in a one-shot setting on Heat.
  1. 1.
  2. 2.
  3. 3.
When modeling uncertainty, the model achieves a traedoff between the quality of the mean prediction approximation and the guarantee for this prediction to be in the corresponding confidence interval. Figure 5 illustrates the trade-off between mean prediction accuracy and uncertainty calibration. At lower temperatures, we achieve the most accurate predictions, but with lower variance, i.e. with no guarantee that the target value is within the confidence interval around the predicted mean. Across most datasets, the confidence level then remains low (less than 80% for $\tau<0.25$), indicating that the true solutions are not reliably captured within the empirical confidence intervals. Conversely, increasing the temperature results in less accurate mean predictions and higher relative standard deviations, but the confidence intervals become more reliable, with levels exceeding 95% for $\tau>0.5$. Therefore, the temperature can be calibrated depending on whether the focus is on accurate point estimates or reliable uncertainty bounds.
Figure 5: Uncertainty quantification with Zebra
5 Limitations
The quality of the generated trajectories is limited by the decoder’s ability to reconstruct details from the quantized latent space. While the reconstructions are excellent for many applications, we believe there is room for improvement. Future work could explore scaling the codebook size, as suggested by Yu et al. (2023a) and Mentzer et al. (2023), to enhance the model’s reconstruction capabilities. Additionally, investigating approaches that avoid vector quantization (Li et al., 2024) could offer even further improvements, provided that in-context learning capabilities are preserved. Lastly, our encoder and decoder are built using convolutional blocks, which restricts their use to regular domains. More flexible architectures, such as those proposed by Serrano et al. (2024), could help extend the model to more complex and irregularly sampled systems.
6 Conclusion
7 Reproducibility Statement
We describe the pretraining strategy in Section 3, and provide details on the architecture and its hyperparameters in Appendix C. The datasets used are described in Appendix B. We plan to release the code, the weights of the models, and the datasets used in this study.
Appendix A Related Work
A.1 Learning parametric PDEs
The classical ML paradigm
The classical ML paradigm for solving parametric PDEs consists in sampling from the PDE parameter distribution trajectories to generalize to new PDE parameter values. It is the classical ERM approach. The natural way for generalizing to new PDE parameters is to explicitly embedded them in the neural network (Brandstetter et al., 2022b). Takamoto et al. (2023) proposed a channel-attention mechanism to guide neural solvers with the physical coefficients given as input; it requires complete knowledge of the physical system and are not designed for other PDE parameter values, e.g., boundary conditions. It is commonly assumed that prior knowledge are not available, but instead rely on past states of trajectories for inferring the dynamics. Neural solvers and operators learn parametric PDEs by stacking the past states as channel information as done in Li et al. (2021), or by creating additional temporal dimension as done in video prediction contexts (Ho et al., 2022; McCabe et al., 2023). Their performance drops when shifts occur in the data distribution, which is often met with parametric PDEs, as small changes in the PDE parameters can lead to various dynamics. To better generalize to new PDE parameter values, Subramanian et al. (2023) instead leverages fine-tuning from pretrained models to generalize to new PDE parameters. It however often necessitates a relatively large number of fine tuning samples to effectively adapt to new PDE parameter values, by updating all or a subset of parameters (Herde et al., 2024; Hao et al., 2024).
Adaptive conditioning
To better adapt to new PDE parameters values at inference, several works have explored meta-learning (also referred to multi-environments learning). During training, a limited number of environments are available, each corresponding to a specific PDE instance. Yin et al. (2022) proposes LEADS, a multi-task framework for learning parametric PDEs, where a shared model from all environments and a model specific to each environment are learned jointly. At inference, for a new PDE instance, the shared model remain frozen and only a model specific to that environment is learned. Kirchmeyer et al. (2022) proposes to perform adaptive conditioning in the parameter space; the framework adapts the weights of a model to each environment via a hyper-network conditioned by a context vector $c^{e}$ specific to each environment. At inference, the model adapts to a new environment by only tuning $c^{e}$. Park et al. (2023) bridges the gap from the classical gradient-based meta-learning approaches by addressing the limitations of second-order optimization of MAML and its variants (Finn et al., 2017; Zintgraf et al., 2019b).
A.2 Generative models
Auto-regressive Transformers for Images and Videos
Recent works have explored combining language modeling techniques with image and video generation, typically using a VQ-VAE (Oord et al., 2017) paired with a causal transformer (Esser et al., 2021) or a bidirectional transformer (Chang et al., 2022). VQGAN (Esser et al., 2021) has become the leading framework by incorporating perceptual and adversarial losses to improve the visual realism of decoder outputs from quantized latent representations. However, while these methods succeed in generating visually plausible images, they introduce a bias—driven by perceptual and adversarial losses—that leads the network to prioritize perceptual similarity and realism, often causing reconstructions to deviate from the true input. In contrast, Zebra focuses on maximizing reconstruction accuracy, and did not observe benefits from using adversarial or perceptual losses during training.
Appendix B Dataset details
@@ B.1 Advection
We fix the size of the domain $L=128$ and draw initial conditions as described in Equation 5 in section B.5 and generate solutions with the method of lines and the pseudo-spectral solver described in Brandstetter et al. (2022b). We take 140 snapshots along a 100s long simulations, which we downsample to 14 timestamps for training. We used a spatial resolution of 256.
@@ B.2 Burgers
We consider the Burgers equation as a special case of the combined equation described in Section B.5 and initially in Brandstetter et al. (2022b), with fixed $\gamma=0$ and $\alpha=0.5$. However, in this setting, we include a forcing term $\delta(t,x)=\sum_{j=1}^{J}A_{j}\sin(\omega_{j}t+2\pi\ell_{j}x/L+\phi_{j})$ that can vary across different environments. We fix $J=5$, $L=16$. We draw initial conditions as described in Equation 5.
@@
We use the solver from Brandstetter et al. (2022b), and take 250 snapshots along the 4s of the generations. We employ a spatial resolution of 256 and downsample the temporal resolution to 25 frames.
@@ B.3 Heat
We consider the heat equation as a special case of the combined equation described in Section B.5 and initially in Brandstetter et al. (2022b), with fixed $\gamma=0$ and $\alpha=0$. However, in this setting, we include a forcing term $\delta(t,x)=\sum_{j=1}^{J}A_{j}\sin(\omega_{j}t+2\pi\ell_{j}x/L+\phi_{j})$ that can vary across different environments. We fix $J=5$, $L=16$. We draw initial conditions as described in Equation 5.
@@
We use the solver from Brandstetter et al. (2022b), and take 250 snapshots along the 4s of the generations. We employ a spatial resolution of 256 and downsample the temporal resolution to 25 frames.
@@ B.4 Wave boundary
We consider a 1D wave equation as in Brandstetter et al. (2022b).
@@
The initial condition is a Gaussian pulse with a peak at a random location. Numerical ground truth is generated with the solver proposed in Brandstetter et al. (2022b). We obtain ground truth trajectories with resolution $(n_{x},n_{t})=(256,250)$, and downsample the temporal resolution to obtain trajectories of shape $(256,60)$.
@@ B.5 Combined equation
We used the setting introduced in Brandstetter et al. (2022b), but with the exception that we do not include a forcing term. The combined equation is thus described by the following PDE:
@@
For training, we sampled 1200 triplets of parameters uniformly within the ranges $\alpha\in[0,1]$, $\beta\in[0,0.4]$, and $\gamma\in[0,1]$. For each parameter instance, we sample 10 trajectories, resulting in 12000 trajectories for training and 120 trajectories for testing. We used the solver proposed in Brandstetter et al. (2022a) to generate the solutions. The trajectories were generated with a spatial resolution of 256 for 10 seconds, along which 140 snapshots are taken. We downsample the temporal resolution to obtain trajectories with shape (256, 14).
@@ C.1 Baseline implementations
  • CODA: For CODA, we implemented a U-Net Ronneberger et al. (2015) and a FNO (Li et al., 2020) as the neural network decoder. For all the different experiments, we reported in the results the best score among the two backbones used. We trained the different models in the same manner as Zebra, i.e. via teacher forcing (Radford et al., 2018). The model is adapted to each environment using a context vector specific to each environment. For the size of the context vector, we followed the authors recommendation and chose a context size equals to the number of degrees of freedom used to define each environment for each dataset. At inference, we adapt to a new environment using 250 gradient steps.
@@
  • [CLS] ViT: For the ViT, we use a simple vision transformer architecture Dosovitskiy et al. (2021), but adapt it to a meta-learning setting where the CLS token encodes the specific variations of each environment. At inference, the CLS token is adapted to a new environment with 100 gradient steps.
  • MPP: For MPP, we used the same model as the one used in the paper (McCabe et al., 2023). As MPP was initially designed for 2D data, we also implemented a 1D version of MPP, to evaluate it both on our 1D and 2D datasets. At inference, MPP can be directly evaluated on new trajectories.
C.2 Zebra additional generation details
We provide illustrations of our inference pipeline in Figure 6 and in Figure 7 both in the case of adaptive conditioning and temporal conditioning. We also include a schematic view of the different generation possibilities with Zebra in Figure 8, using the sequence design adopted during pretraining
@@ C.2 Zebra additional generation details
Figure 7: Zebra’s inference pipeline from observations of several initial frames. The initial timestamps are tokenized into index sequences that are concatenated according to the sequence design adopted during pretraining. The transformer then generates the next tokens to complete the sequence. We detokenize these indices to get back to the physical space.

### v3

@@ B.1 Advection
We fix the size of the domain $L=128$ and draw initial conditions as described in Equation 3 in section B.5 and generate solutions with the method of lines and the pseudo-spectral solver described in (Brandstetter et al., 2022b). We take 140 snapshots along a 100s long simulations, which we downsample to 14 timestamps for training. We used a spatial resolution of 256.
@@ B.2 Burgers
We consider the Burgers equation as a special case of the combined equation described in Section B.5 and initially in (Brandstetter et al., 2022b), with fixed $\gamma=0$ and $\alpha=0.5$. However, in this setting, we include a forcing term $\delta(t,x)=\sum_{j=1}^{J}A_{j}\sin(\omega_{j}t+2\pi\ell_{j}x/L+\phi_{j})$ that can vary across different environments. We fix $J=5$, $L=16$. We draw initial conditions as described in Equation 3.
@@
We use the solver from (Brandstetter et al., 2022b), and take 250 snapshots along the 4s of the generations. We employ a spatial resolution of 256 and downsample the temporal resolution to 25 frames.
@@ B.3 Heat
We consider the heat equation as a special case of the combined equation described in Section B.5 and initially in (Brandstetter et al., 2022b), with fixed $\gamma=0$ and $\alpha=0$. However, in this setting, we include a forcing term $\delta(t,x)=\sum_{j=1}^{J}A_{j}\sin(\omega_{j}t+2\pi\ell_{j}x/L+\phi_{j})$ that can vary across different environments. We fix $J=5$, $L=16$. We draw initial conditions as described in Equation 3.
@@
We use the solver from (Brandstetter et al., 2022b), and take 250 snapshots along the 4s of the generations. We employ a spatial resolution of 256 and downsample the temporal resolution to 25 frames.
@@ B.4 Wave boundary
We consider a 1D wave equation as in (Brandstetter et al., 2022b).
@@
The initial condition is a Gaussian pulse with a peak at a random location. Numerical ground truth is generated with the solver proposed in (Brandstetter et al., 2022b). We obtain ground truth trajectories with resolution $(n_{x},n_{t})=(256,250)$, and downsample the temporal resolution to obtain trajectories of shape $(256,60)$.
@@ B.5 Combined equation
We used the setting introduced in (Brandstetter et al., 2022b), but with the exception that we do not include a forcing term. The combined equation is thus described by the following PDE:
@@
For training, we sampled 1200 triplets of parameters uniformly within the ranges $\alpha\in[0,1]$, $\beta\in[0,0.4]$, and $\gamma\in[0,1]$. For each parameter instance, we sample 10 trajectories, resulting in 12000 trajectories for training and 120 trajectories for testing. We used the solver proposed in (Brandstetter et al., 2022a) to generate the solutions. The trajectories were generated with a spatial resolution of 256 for 10 seconds, along which 140 snapshots are taken. We downsample the temporal resolution to obtain trajectories with shape (256, 14).
@@ C.1 Baseline implementations
  • CODA: For CODA, we implemented a U-Net (Ronneberger et al., 2015) and a FNO (Li et al., 2020) as the neural network decoder. For all the different experiments, we reported in the results the best score among the two backbones used. We trained the different models in the same manner as Zebra, i.e. via teacher forcing (Radford et al., 2018). The model is adapted to each environment using a context vector specific to each environment. For the size of the context vector, we followed the authors recommendation and chose a context size equals to the number of degrees of freedom used to define each environment for each dataset. At inference, we adapt to a new environment using 250 gradient steps.
@@
  • [CLS] ViT: For the ViT, we use a simple vision transformer architecture (Dosovitskiy et al., 2021), but adapt it to a meta-learning setting where the CLS token encodes the specific variations of each environment. At inference, the CLS token is adapted to a new environment with 100 gradient steps.
  • ViT-in-context: We implement ViT-in-context using a standard transformer architecture with separate temporal and spatial attention mechanisms, following Ho et al. (2019). During both training and inference, context examples are stacked along the temporal dimension. The model is trained to predict the next frame in the target sequence, conditioned on both the context examples and the preceding frames of the target sequence.
C.2 Zebra additional details
Zebra
We describe the pretraining strategy in Section 3, and provide details on the architecture and its hyperparameters in Appendix C. The datasets used are described in Table 6. We plan to release the code, the weights of the models, and the datasets used in this study upon acceptance.
For clarity, we outline the pretraining steps of Zebra in Figure 5 illustrated with the vorticity 2D dataset.
Figure 5: Zebra’s pretraining includes two stages. 1) A finite vocabulary of physical phenomena is learned by training a VQ-VAE on spatial representations. 2) During the pretraining, multiple trajectories sharing the same dynamics are tokenized and concatenated into a common sequence $S$. The transformer is trained to predict the next-token by minimizing the cross-entropy loss.
We also provide illustrations of our inference pipeline in Figure 6. We finally include a schematic view of the different generation possibilities with Zebra in Figure 8, using the sequence design adopted during pretraining.
@@ C.2 Zebra additional details
Zebra + UNet
Zebra is competitive both in-distribution and out-of-distribution, while also enabling uncertainty quantification due to its generative nature. Additionally, Zebra can generate novel trajectories and initial conditions, providing a way to sample complex initial states. As such Zebra is already faster than gradient-based adaptation methods, but since the model generates trajectory solutions token by token, the number of calls to the transformer increases by one or two orders of magnitude compared to direct surrogate modeling, making the process costly.
In this section, we propose a hybrid approach that leverages Zebra’s pretrained knowledge in combination with a conventional neural surrogate model. The objective is to develop a framework that encodes context trajectories into an embedding vector, which then conditions a neural surrogate, similar to classical adaptation methods such as CODA and CAPE.
To achieve this, we finetune Zebra as an encoder to adapt a conditional surrogate model, such as a UNet (Ronneberger et al., 2015). We introduce a [DYN] token (short for dynamics), which is appended to the right of the context sequence during both training and inference. This allows Zebra to extract a dynamics embedding from the transformer’s output, defined as:
[…]
where $S$ represents a sequence of tokens encoding the context trajectories. The embedding $\xi_{S}$ captures key properties of the dynamics and is mapped to the conditioning space of a UNet, which adapts the model to each specific dynamics. Following Gupta & Brandstetter (2022), the UNet conditioning modifies the network’s biases. Once the dynamics embedding is extracted, the UNet can directly predict the next state from the current state :
$\hat{{\bm{u}}}^{t+\Delta t}=\text{UNet}({\bm{u}}^{t},\xi_{S}).$
With this architecture, we effectively extract the key dynamics from example trajectories and use this information to adapt a neural surrogate, significantly accelerating inference. The complete pipeline at inference is illustrated in Figure 7.
To efficiently finetune Zebra, we apply LoRA (Hu et al., 2021), keeping the transformer’s weights frozen while learning low-rank updates. This setup enables the UNet to leverage Zebra’s pre-learned representations while achieving a substantial speedup. Compared to standalone Zebra, integrating a UNet improves inference speed by a factor of ×$\times 30$ in 1D and ×$\times 150$ in 2D. The method is also considerably faster than CODA and CAPE while maintaining competitive performance across tasks. The inference times are summarized in Table 4.
Finally we illustrate the inference pipeline for accelerating the inference of Zebra with the UNet in Figure 7.
Figure 7: Zebra + UNet inference pipeline. Zebra serves as an encoder, utilizing the special token [DYN] to generate a dynamics embedding $\xi$ from the context example trajectories. Once this embedding is obtained, the UNet can autoregressively forecast the sequence significantly faster than a next-token transformer.
C.3 Auto-regressive transformer